1) Identify the fifth term in the sequence in which a^2=40 and r=1/2
2) Find the sum of the geometric series in which a^1=125,r=2/5, and n=4

Answer by venugopalramana(3286)

1) Identify the fifth term in the sequence in which a^2=40 and r=1/2
TN=A*R^(N-1)
T5=A*R^(5-1)=A*R^4=SQRT(40)*(1/2)^4=SQRT(40)/16=2SQRT(10)/16=SQRT(10)/8=0.3953

2) Find the sum of the geometric series in which a^1=125,r=2/5, and n=4
sn=a*(1-r^n)/(1-r)
S4=125*(1-(2/5)^4)/(1-(2/5))=203